A hole in a brass plate has a diameter of 20mm through what temperature rise must the plate be heated for the diameter to become 20.09mm
1
Expert's answer
2019-07-25T17:09:40-0400
"d'=d(1+\alpha \Delta T)"

"\Delta T=\frac{(\frac{d'}{d}-1)}{\alpha}"

"\Delta T=\frac{(\frac{20.09}{20}-1)}{19\cdot 10^{-6}}=240\ K"